OpenStudy (anonymous):

What is 36x^2-51x+18 factored?

OpenStudy (campbell_st):

start by factoring out 3 \[3[12x^2 - 17x + 6]\] so looking at the quadratic... a never fail method for \[ax^2 + bx + c\] is to multiply a and c so you have \[12 \times 6 = 72\] find the factors of 72 that add to -17... both factors are negative. any thoughts on the factors...?

OpenStudy (anonymous):

3(3x−2)(4x−3)

OpenStudy (campbell_st):

there you go, someone always gives the answer... why not show your working...

OpenStudy (anonymous):

Factors are -9 &-8

OpenStudy (campbell_st):

so then the expression is written as (ax + factor 1)(ax + factor 2) ------------------------- a so your problem becomes \[3 \times \frac{(12x -9)(12x -8)}{12}\] now remove the common factors from the binomials \[3 \times \frac{4(3x -2)3(4x - 3)}{12}\] then cancel the common factors leaving the answer posted above.. works every time a difficult quadratic can be factored...